Interim Constitution of the SHIP Collaboration

(proposed 14 February 2014) (modified 18 February 2014) (adopted 20 February 2014) (modified and adopted 18 December 2014)

I - Preamble

* The aim of the SHIP (Search for HIdden Particles) Collaboration is to prepare a proposal for an experiment to be performed at CERN, with as main goal the physics described in CERN-SPSC-2013-024 / SPSC-EOI-010 (http://arxiv.org/abs/1310.1762), and to perform the experiment when approved.

* The term “proposal” is used in this constitution to mean any document that is required by CERN to approve the experiment, such as a LOI, proposal, CDR, or TDR.

* The interim Constitution of the SHIP Collaboration describes the management of the Collaboration during the time that the work concentrates on producing a proposal to be submitted to CERN.

* The interim Constitution is meant to be updated as soon as the situation requires a more elaborate structure.

* In any case, the interim Constitution will be replaced by a new Constitution when the proposal is accepted by CERN.

* During this initial phase, an interim Steering Group (iSG) is the policy making and decision making body of the Collaboration.

* An interim Spokesperson (iSP) is responsible for external contacts and contacts with the CERN management.

* The iSG leads the operation of the Collaboration and can delegate the daily tasks to the iSP.

* Membership issues are decided by the Country Representatives Board (CRB).

* Working Groups lead by Convenors can be set up by the iSG to facilitate the progress of the project.

II - Interim Steering Group

* The iSG is the policy making and decision making body of the Collaboration.

* Membership and voting rights

The iSG is composed of those who have signed the original EOI and can co-opt new members when deemed necessary to perform its tasks. The Members of the iSG are listed in Appendix A. The iSG is chaired by the iSP.

The iSG has frequent meetings, typically once per week, to ensure the progress of the work.

* Decision Procedures

The decisions in the iSG are taken by consensus whenever possible or otherwise by vote. Each member represented in the iSG has one vote. All votes are open except for elections and re-elections. Decisions of the iSG require a 50% majority of the valid votes. Abstentions, invalid and null votes are not taken into account for the definition of a majority. Thus in case of a yes/no vote only the sum of the positive and negative votes will be the 100% value. In case of an election the sum of valid votes for anyone of the candidates will be the 100% value. A change of the Constitution needs a 2/3 majority. Elections are performed by a succession of voting rounds whereby the candidate who has received more than 50% of the votes is elected. If no candidate receives this majority, the candidate(s) with the smallest number of votes is (are) removed in a further election round. The process is repeated until a majority is reached. In case no majority is reached, or when all candidates would be removed from the next election round by the rule of least votes, the candidate with the highest age is removed from the next iteration.

* Country Representatives Board (CRB)

The iSG sets up the Country Representatives Board, which has as task to consider and prepare the adoption of new members of the Collaboration following rules defined in this document. Each country (with members in the Collaboration) has one representative. In case there is more than one Funding Agency per country, there is one representative per Funding Agency.

* Working Groups and Convenors

The iSG sets up working groups as deemed necessary to advance the work for the proposal. The iSG, nominates the Convenors of the Working Groups with an election. The Convenors report to the iSG. The iSG delegates the daily interaction with the Working Groups to the iSP.

* Agenda and Minutes

The agenda for the iSG is prepared by its chairperson, and is circulated in advance. Any member of the iSG may ask for additional topics to be included. The proceedings of the iSG are recorded in minutes. Draft minutes are circulated to members of the iSG for approval at the next meeting. All iSG members will be informed at least one day in advance about all subjects on which a decision must be taken.

III - Membership and Associated Members of the Collaboration and Country Representatives Board (CRB)

* Membership rules

Membership rules distinguish between personal members (“Members”) and Institute Members.

* Membership of the Collaboration

Members of the Collaboration are those who have signed the original EOI or contributed to the replies to the questions of the SPSC referees. Additional members are accepted according to the rules laid out for “New Members”. The list of Members is maintained as Appendix B.

* Institute Membership of the Collaboration

Member Institutes of the Collaboration are all those institutes to which the original members are affiliated. New institutes can be accepted as Member Institutes following rules laid down in this document. The list of Member and Associated Member Institutes is maintained as Appendix B. Member Institutes appoint a person who acts as the contact with the iSG and iSP.

* New Members and new Member Institutes of the Collaboration

The candidature of new Member Institutes and new Members is first considered by the Chairman of the Country Representative Board (CRB) and the Spokesperson. Once they decide that sufficient information is available on the intended participation, the candidature is presented to the (CRB) for a vote.

* Associated Members of the Collaboration

An Institute can be given the status of Associated Member if it wishes to bring in limited technical, financial or human resources to a particular project or working group.

An Associated Member must be ”attached” to an Institute of the SHiP Collaboration, called the ”Host Institute”. The Host Institute will propose the Associated Member candidature to the Country Representative Board.

The Host Institute must take the long-term responsibility for the operation and maintenance of any hardware, or software, which the Associated Member would produce and which would become part of the experiment.

* Country Representative Board (CRB)

The composition and chairperson of the CRB is elected by the iSG, following the procedure defined for all iSG elections. The composition of the CRB reflects the needs for a regional representation and representation of the physics community for the adoption of new members. The composition of the CRB is maintained as Appendix C. The CRB prepares membership decisions to be taken in the iSG. The CRB operates by consensus whenever possible or otherwise by vote. In case a vote is needed, the CRB follows the same rules as the iSG.

IV - Management Functions

* General Rules

All mandates and appointments are valid for a limited period of time as defined by the iSG. More than one function may be held simultaneously with the agreement of the iSG. All post holders may be re-elected.

* Interim Spokesperson (iSP)

The iSP is responsible to the iSG for the execution of the project. The iSP represents the Collaboration to the SPSC, to CERN Management, and to the outside world. The iSP is elected by the iSG, following the procedure defined for all iSG elections. The election of the iSP is organised by a small group of persons appointed by the iSG. Candidates should be presented by at least one Member of the iSG and their names should be announced at least two days before the election. The iSG may decide at any time with a 50% majority to start a new election procedure of the interim Spokesperson. A new Spokesperson election will in any case be started when the proposal is approved by CERN.

* Convenors

The Convenors report to the iSG and are responsible for the operation of their Working Group. Convenors are elected by the iSG, following the procedure defined for all iSG elections. The iSG may decide at any time to replace a Convenor, or set up a new Working Group.